OpenStudy (anonymous):

last question: negative 2 third times 3 fourths = A. -1/2 B. -9/8 C. 1/2 D. 9/8 @kenziejo1998

OpenStudy (anonymous):

well when you multiply -2/3 by 3/4 you get -.5 which is??

OpenStudy (anonymous):

@roger19garcia

OpenStudy (anonymous):

i dont quite understand

OpenStudy (anonymous):

what is -.5 in a fraction??

OpenStudy (anonymous):

you know it just try. you can do it

OpenStudy (anonymous):

your answer would be A, -1/2

OpenStudy (anonymous):

i was trying to get her to answer without giving just the answer....... LOL oh well. but the answer is A

OpenStudy (anonymous):

-2 x 3 = -6 3 x 4 = 12 -6/12 = -1/2

OpenStudy (anonymous):

thanks guys im so glad to have your help

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

yw